• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 21:55:08 on March 10, 2009

    Entry number 265571

    This entry is a followup to: 265560

    Followups:

    keyword=Compton line pump-down update . . . beamline valves closed until tomorrow?

    ELOG 1461653 from 18:42 is reproduced here:
    The compton electron detector in hall A was let up to facilitate maintanance per Sirish Nanda.After the thin metal foil was installed in the detector a leak test was performed and no leaks were found. The pressure in the detector beam line is 2.7 e-05 at the turbo pump and coming down very slowly. A decision was made to pump with the turbo pump over night and try to start the ion pumps in the morning.The valves will also be re-connected in the morning. We are hoping that the material used for the shielding is not out gassing.

    Figure 1 is shows the pressure vs. time in the compton line.

    A copy of this log entry has been emailed to: riordan